OpenStudy (anonymous):

I need help ! Sole the equation for the exact values using the square root principle. (2x-1)²+4=0

OpenStudy (ddcamp):

Square root principle: \[\sqrt{a^2} = a\] So we can say: \[(2x-1)^2+4=0\\(2x-1)^2=-4\\\sqrt{(2x-1)^2}=\sqrt{-4}\\2x-1=\sqrt{-4}\\2x=\sqrt{-4}+1\\x=\frac{\sqrt{-4}+1}{2}\]

OpenStudy (anonymous):

thank you so much